OpenStudy (anonymous):

The position of a 2kg mass is given by x=(2t^3-3t^2)m, where t is in seconds. What is the net horizontal force on the mass at (a)t=0s and (b)t=1s?

OpenStudy (mos1635):

x=(2t^3-3t^2 v=dx/dt v=6t^2-6t a=dv/dt a= 12t-6 F=m*a F=2*(12t-6) F=24t-12 t=0 => F=-12 t=1 => F=12